# Ring Attention with Blockwise Transformers for Near-Infinite Context

### v1

@@ Abstract
Transformers have emerged as the architecture of choice for many state-of-the-art AI models, showcasing exceptional performance across a wide range of AI applications. However, the memory demands imposed by Transformers limit their ability to handle long sequences, thereby creating challenges for tasks involving extended sequences or long-term dependencies. We present a distinct approach, Ring Attention, which leverages blockwise computation of self-attention to distribute long sequences across multiple devices while concurrently overlapping the communication of key-value blocks with the computation of blockwise attention. By processing longer input sequences while maintaining memory efficiency, Ring Attention enables training and inference of sequences that are device count times longer than those of prior memory-efficient Transformers, effectively eliminating the memory constraints imposed by individual devices. Extensive experiments on language modeling tasks demonstrate the effectiveness of Ring Attention in allowing large sequence input size and improving performance.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have become the backbone of many state-of-the-art AI systems that have demonstrated impressive performance across a wide range of AI problems. Transformers achieve this success through their architecture design that uses self-attention and position-wise feedforward mechanisms.
Figure 1: Maximum context length on TPUv4-512 (32GB memory on each TPUv4). Baselines are vanilla transformers (Vaswani et al., 2017), transformers with memory efficient attention (Rabe and Staats, 2021), and memory efficient attention and feedforward (blockwise parallel transformers) (Liu and Abbeel, 2023b). Our proposed approach Ring Attention allows training 512 times longer sequence than baselines and enables the training of sequences that exceed 100 million in length without making approximations to attention.
These components facilitate the efficient capture of long-range dependencies between input tokens, and enable scalability through highly parallel computations.
@@
However, scaling up the context length of Transformers is a challenge (OpenAI, 2023), since the inherited architecture design of Transformers, i.e. the self-attention has memory cost quadratic in the input sequence length, which makes it challenging to scale to longer input sequences. Large context Transformers are essential for tackling a diverse array of AI challenges, ranging from processing books and high-resolution images to analyzing long videos and complex codebases. They excel at extracting information from the interconnected web and hyperlinked content, and are crucial for handling complex scientific experiment data. There have been emerging use cases of language models with significantly expanded context than before: GPT-3.5 (Schulman et al., 2022) with context length 16K, GPT-4 (OpenAI, 2023) with context length 32k, MosaicML’s MPT (MosaicML, 2023) with context length 65k, and Anthropic’s Claude (Anthropic, 2023) with context length 100k.
@@
Driven by the significance, there has been surging research interests in reducing memory cost. One line of research leverages the observation that the softmax matrix in self-attention can be computed without materializing the full matrix (Milakov and Gimelshein, 2018) which has led to the development of blockwise computation of self-attention and feedforward (Rabe and Staats, 2021; Dao et al., 2022; Liu and Abbeel, 2023b) without making approximations. Despite the reduced memory, a significant challenge still arises from storing the output of each layer. This necessity arises from self-attention’s inherent nature, involving interactions among all elements (n to n interactions). The subsequent layer’s self-attention relies on accessing all of the prior layer’s outputs. Failing to do so would increase computational costs cubically, as every output must be recomputed for each sequence element, rendering it impractical for longer sequences. To put the memory demand in perspective, even when dealing with a batch size of 1, processing 100 million tokens requires over 1000GB of memory for a modest model with a hidden size of 1024. This is much greater than the capacity of contemporary GPUs and TPUs, which typically have less than 100GB of high-bandwidth memory (HBM).
@@
To tackle this challenge, we make a key observation: by performing self-attention and feedforward network computations in a blockwise fashion (Liu and Abbeel, 2023b), we can distribute sequence dimensions across multiple devices, allowing concurrent computation and communication. This insight stems from the fact that when we compute the attention on a block-by-block basis, the results are invariant to the ordering of these blockwise computations. Our method distributes the outer loop of computing blockwise attention among hosts, with each device managing its respective input block. For the inner loop, every device computes blockwise attention and feedforward operations specific to its designated input block. Host devices form a conceptual ring, where during the inner loop, each device sends a copy of its key-value blocks being used for blockwise computation to the next device in the ring, while simultaneously receiving key-value blocks from the previous one. Because block computations take longer than block transfers, overlapping these processes results in no added overhead compared to standard transformers. By doing so, each device requires memory only proportional to the block size, which is independent of the original input sequence length. This effectively eliminates the memory constraints imposed by individual devices. Since our approach overlaps the communication of key-value blocks between hosts in a ring with blockwise computation, we name it Ring Attention.
@@ 2 Large Context Memory Constraint
Large Output of Each Layer. While BPT significantly reduces memory demand in Transformers, it still presents a major challenge for scaling up context length because it requires storing the output of each layer. This storage is crucial due to the inherent nature of self-attention, which involves interactions among all elements (n to n interactions). Without these stored outputs, the subsequent layer’s self-attention becomes computationally impractical, necessitating recomputation for each sequence element. To put it simply, processing 100 million tokens with a batch size of 1 requires over 1000GB of memory even for a modest model with a hidden size of 1024. In contrast, modern GPUs and TPUs typically provide less than 100GB of high-bandwidth memory (HBM), and the prospects for significant GPU HBM expansion are hindered by physical limitations and high manufacturing costs.
3 Ring Attention
@@
Figure 2: Top (a): We use the same model architecture as the original Transformer but with a different way of organizing the compute. In the diagram, we explain this by showing that in a ring of hosts, each host holds one query block, and key-value blocks traverse through hosts for attention and feedforward computations in a block-by-block fashion. As we compute attention, each host sends key-value blocks to the next host while receives key-value blocks from the preceding host. The communication is overlapped with the computation of blockwise attention and feedforward. Bottom (b): We compute the original Transformer block-by-block. Each host is responsible for one iteration of the query’s outer loop, while the key-value blocks rotate among the hosts. As visualized, a device starts with the first query block on the left; then we iterate over the key-value blocks sequence positioned horizontally. The query block, combined with the key-value blocks, are used to compute self-attention (yellow box), whose output is pass to feedforward network (cyan box).
@@
Ring-Based Blockwise Attention. To tackle the aforementioned challenges, we leverage the permutation invariance property of the inner loop’s key-value block operations. This property stems from the fact that the self-attention between a query block and a group of key-value blocks can be computed in any order, as long as the statistics of each block are combined correctly for rescaling. We leverage this property by conceptualizing all hosts as forming a ring structure: host-$1$, host-$2$, …, host-$N$. As we compute blockwise attention and feedforward, each host efficiently coordinates by concurrently sending key-value blocks being used for attention computation to the next host while receiving key-value blocks from the preceding host, effectively overlapping transferring of blocks with blockwise computation. Concretely, for any host-$i$, during the computation of attention between its query block and a key-value block, it concurrently sends key-value blocks to the next host-$(i+1)$ while receiving key-value blocks from the preceding host-$(i-1)$. If the computation time exceeds the time required for transferring key-value blocks, this results in no additional communication cost. This overlapping mechanism applies to both forward and backward passes of our approach since the same operations and techniques can be used.
@@
Our analysis shows that the model needs to fit in $s=6c$ sequence length, i.e., six times of minimal block size. Requirements on popular computing servers as shown in Table 2, the required minimal sequence length to be fit in each host is between 6K to 20K. This requirement is easy to meet using memory efficient blockwise attention and feedforward (Rabe and Staats, 2021; Dao et al., 2022; Liu and Abbeel, 2023b), which we will show in experiment section 5.
@@ 5.1 Evaluating Max Context Size
We evaluate maximum supported context length using tensor parallelism and batch size 1 in sequences. Following prior works (Liu and Abbeel, 2023b; Shoeybi et al., 2019), we note that no data parallelism is considered in our evaluations since our approach is independent of data parallelism. As a result, the batch sizes used in our analysis are much lower than the ones used for the end-to-end training. Practitioners can combine our method with data parallelism to scale up batch size, which we will show in Section 5.2. Table 3 summarizes the results of our experiments.
Our Ring Attention model consistently surpasses baselines, delivering superior scalability across diverse hardware setups. For example, with 32 A100 GPUs, we achieve over 32 million tokens in context size, a significant improvement over baselines. Furthermore, when utilizing larger accelerators like TPUv4-512, Ring Attention enables a 512x increase in context size, allows training sequences of over 100 million tokens. Furthermore, our Ring Attention model scales linearly with the number of devices, as demonstrated by the 8x improvement over BPT on 8 A100 and the 512x improvement on TPUv4-512. If a model can be trained with context size $s$ on $n$ GPUs using the blockwise attention and feedforward, with our Ring Attention approach, it becomes possible to train a model with a context size of $ns$.
Table 3: Maximum context length supported in device memory on different model sizes and clusters of accelerators. Baselines are vanilla transformer (Vaswani et al., 2017), transformer with memory efficient attention (Rabe and Staats, 2021), and transformer with memory efficient attention and feedforward (Liu and Abbeel, 2023b). The context size is reported in tokens (1e3). Our Ring Attention substantially outperforms baselines and enables training sequences that are device count times longer than prior state-of-the-arts, achieving over 100M context size.
@@
7B 32 512 1024 32768 (32M) 32x
30B 16 128 512 16384 (16M) 32x
TPUv3-512  [^1] Unlike TPUv4-256 and TPUv5-256 where the number 256 represents the count of TPUv4 (v5) hosts, TPUv3 uses a doubled host count notation. So, TPUv3-512 means there are 256 hosts. See https://cloud.google.com/tpu/docs/system-architecture-tpu-vm#tpu_v3 for more details.
7B 4 16 64 16384 (16M) 256x
13B 2 8 32 8192 (8M) 256x
30B 1 4 16 4096 (4M) 256x
TPUv4-512
3B 8 64 256 131072 (131M) 512x
7B 8 32 128 65536 (65M) 512x
13B 4 16 64 32768 (32M) 512x
30B 2 8 32 16384 (16M) 512x
@@
7B 4 16 64 16384 (16M) 256x
30B 1 4 16 4096 (4M) 256x
Table 4: Model flops utilization (MFU) with different training configurations: model sizes, compute, and context lengths. Ring Attention enables training large models (30B-65B) on large input context sizes (over 8M) with negligible overheads.
Model size 7B 13B 30B 30B 65B
Compute 8x A100 8x A100 32x A100 TPUv4-256 TPUv4-256
Memory efficient attention & FFN Context size (×$\times 1\mathrm{e}3$) 64 32 64 32 16
Ring Attention Context size (×$\times 1\mathrm{e}3$) 512 256 512 8192 (8M) 4096 (4M)
@@ 5.2 Evaluating Model Flops Utilization
We evaluate the model flops utilization (MFU) of Ring Attention in standard training settings using fully sharded data parallelism(FSDP) (Facebook, 2023) and tensor parallelism following LLaMA and OpenLLaMA (Touvron et al., 2023; Geng and Liu, 2023). The batch size in tokens are 2M on 8/32x A100 and 4M on TPUv4-256. Our goal is investigating the impact of model size and context length on MFU, a critical performance metrics while highlighting the benefits of our approach. Table 4 presents the results of our experiments on MFU for different model sizes and context lengths. We present the achieved MFU using state-of-the-art memory efficient transformers BPT (Liu and Abbeel, 2023b), compare it to our anticipated MFU based on these results, and demonstrate the actual MFU obtained with our approach (Ring Attention). For fair comparison, both BPT and our approach are based on the same BPT implementation on both GPUs and TPUs. It’s worth noting that on GPUs our approach Ring Attention can be also integrated with the more compute efficient Triton code (kernel team, 2023) or CUDA code (Dao et al., 2022) of memory efficient attention (Rabe and Staats, 2021), similarly on TPUs it is also compatible with Pallas (team, 2023). Combing these low level kernels implementations with our approach can maximize MFU, we leave that to future work.

### v4

@@ Abstract
Transformers have emerged as the architecture of choice for many state-of-the-art AI models, showcasing exceptional performance across a wide range of AI applications. However, the memory demands imposed by Transformers limit their ability to handle long sequences, thereby posing challenges in utilizing videos, actions, and other long-form sequences and modalities in complex environments. We present a novel approach, Ring Attention with Blockwise Transformers (Ring Attention), which leverages blockwise computation of self-attention and feedforward to distribute long sequences across multiple devices while fully overlapping the communication of key-value blocks with the computation of blockwise attention. Our approach enables training and inference of sequences that are up to device count times longer than those achievable by prior memory-efficient Transformers, without resorting to approximations or incurring additional communication and computation overheads. Extensive experiments on language modeling and reinforcement learning tasks demonstrate the effectiveness of our approach in allowing millions of tokens context size and improving performance.  [^1] Code: https://github.com/lhao499/llm_large_context.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have become the backbone of many state-of-the-art AI systems that have demonstrated impressive performance across a wide range of AI problems. Transformers achieve this success through their architecture design that uses self-attention and position-wise feedforward mechanisms. However, scaling up the context length of Transformers is a challenge (OpenAI, 2023), since the inherited architecture design of Transformers, i.e. the self-attention has memory cost quadratic in the input sequence length, which makes it challenging to scale to longer input sequences. Large context Transformers are essential for tackling a diverse array of AI challenges, ranging from processing books and high-resolution images to analyzing long videos and complex codebases. They excel at extracting information from the interconnected web and hyperlinked content, and are crucial for handling complex scientific experiment data. There have been emerging use cases of language models with significantly expanded context than before: GPT-3.5 (Schulman et al., 2022) with context length 16K, GPT-4 (OpenAI, 2023) with context length 32k, MosaicML’s MPT (MosaicML, 2023) with context length 65k, and Anthropic’s Claude (Anthropic, 2023) with context length 100k.
@@
Driven by the significance, there has been surging research interests in reducing memory cost. One line of research leverages the observation that the softmax matrix in self-attention can be computed without materializing the full matrix (Milakov and Gimelshein, 2018) which has led to the development of blockwise computation of self-attention and feedforward (Rabe and Staats, 2021; Dao et al., 2022; Liu and Abbeel, 2023b) without making approximations. Despite the reduced memory, a significant challenge still arises from storing the output of each layer. This necessity arises from self-attention’s inherent nature, involving interactions among all elements (n to n interactions). The subsequent layer’s self-attention relies on accessing all of the prior layer’s outputs. Failing to do so would increase computational costs cubically, as every output must be recomputed for each sequence element, rendering it impractical for longer sequences.
Figure 1: Maximum context length under end-to-end large-scale training on TPUv4-1024. Baselines are vanilla transformers (Vaswani et al., 2017), memory efficient transformers (Rabe and Staats, 2021), and memory efficient attention and feedforward (blockwise parallel transformers) (Liu and Abbeel, 2023b). Our proposed approach Ring Attention allows training up to device count times longer sequence than baselines and enables the training of sequences that exceed millions in length without making approximations nor adding any overheads to communication and computation.
These components facilitate the efficient capture of long-range dependencies between input tokens, and enable scalability through highly parallel computations. To put the memory demand in perspective, even when dealing with a batch size of 1, processing 100 million tokens requires over 1000GB of memory for a modest model with a hidden size of 1024. This is much greater than the capacity of contemporary GPUs and TPUs, which typically have less than 100GB of high-bandwidth memory (HBM).
@@
To tackle this challenge, we make a key observation: by performing self-attention and feedforward network computations in a blockwise fashion (Liu and Abbeel, 2023b), we can distribute sequence dimensions across multiple devices, allowing concurrent computation and communication. This insight stems from the fact that when we compute the attention on a block-by-block basis, the results are invariant to the ordering of these blockwise computations. Our method distributes the outer loop of computing blockwise attention among hosts, with each device managing its respective input block. For the inner loop, every device computes blockwise attention and feedforward operations specific to its designated input block. Host devices form a conceptual ring, where during the inner loop, each device sends a copy of its key-value blocks being used for blockwise computation to the next device in the ring, while simultaneously receiving key-value blocks from the previous one. As long as block computations take longer than block transfers, overlapping these processes results in no added overhead compared to standard transformers. The use of a ring topology for computing self-attention has also been studied in prior work (Li et al., 2023b) but it incurs non-overlapped communication overheads similar to sequence parallelism, making it infeasible for large context sizes. Our work utilizes blockwise parallel transformers (Liu and Abbeel, 2023b) to substantially reduce memory costs, enabling zero-overhead scaling of context size across tens of millions of tokens during both training and inference, and allowing for the use of an arbitrarily large context size. Since our approach overlaps the communication of key-value blocks between hosts in a ring through blockwise computation of transformers, we name it Ring Attention with Blockwise Parallel Transformers (Ring Attention).
@@ 2 Large Context Memory Constraint
Large Output of Each Layer. While BPT significantly reduces memory demand in Transformers, it still presents a major challenge for scaling up context length because it requires storing the output of each layer. This storage is crucial due to the inherent nature of self-attention, which involves interactions among all elements (n to n interactions). Without these stored outputs, the subsequent layer’s self-attention becomes computationally impractical, necessitating recomputation for each sequence element. To put it simply, processing 100 million tokens with a batch size of 1 requires over 1000GB of memory even for a modest model with a hidden size of 1024. In contrast, modern GPUs and TPUs typically provide less than 100GB of high-bandwidth memory (HBM), and the prospects for significant HBM expansion are hindered by physical limitations and high manufacturing costs.
3 Ring Attention with Blockwise Parallel Transformers
@@
Figure 2: Top (a): We use the same model architecture as the original Transformer but reorganize the compute. In the diagram, we explain this by showing that in a ring of hosts, each host holds one query block, and key-value blocks traverse through a ring of hosts for attention and feedforward computations in a block-by-block fashion. As we compute attention, each host sends key-value blocks to the next host while receives key-value blocks from the preceding host. The communication is overlapped with the computation of blockwise attention and feedforward. Bottom (b): We compute the original Transformer block-by-block. Each host is responsible for one iteration of the query’s outer loop, while the key-value blocks rotate among the hosts. As visualized, a device starts with the first query block on the left; then we iterate over the key-value blocks sequence positioned horizontally. The query block, combined with the key-value blocks, are used to compute self-attention (yellow box), whose output is pass to feedforward network (cyan box).
@@
Ring-Based Blockwise Attention. To tackle the aforementioned challenges, we leverage the permutation invariance property of the inner loop’s key-value block operations. This property stems from the fact that the self-attention between a query block and a group of key-value blocks can be computed in any order, as long as the statistics of each block are combined correctly for rescaling. We leverage this property by conceptualizing all hosts as forming a ring structure: host-$1$, host-$2$, …, host-$N$. As we compute blockwise attention and feedforward, each host efficiently coordinates by concurrently sending key-value blocks being used for attention computation to the next host while receiving key-value blocks from the preceding host, effectively overlapping transferring of blocks with blockwise computation. Concretely, for any host-$i$, during the computation of attention between its query block and a key-value block, it concurrently sends key-value blocks to the next host-$(i+1)$ while receiving key-value blocks from the preceding host-$(i-1)$. If the computation time exceeds the time required for transferring key-value blocks, this results in no additional communication cost. This overlapping mechanism applies to both forward and backward passes of our approach since the same operations and techniques can be used. Prior work has also proposed leveraging a ring topology to compute self-attention (Li et al., 2023b), aiming to reduce communication costs. Our work differs by utilizing blockwise parallel transformers to substantially reduce memory costs. As we show in the next section, this enables zero-overhead scaling of context size during both training and inference and allows arbitrarily large context size.
@@
Our analysis shows that the model needs to have a sequence length of $s=6c$, which is six times the minimal block size. Requirements for popular computing servers are shown in Table 2. The required minimal sequence length (rightmost column) for each host varies between 6K and 10K, and the minimal block size (second-to-rightmost column) for each host is around 1K for TPUs and GPUs with high bandwidth interconnect. For GPUs connected via InfiniBand, which offers lower bandwidth, the requirements are more strict. These requirements are easy to meet with parallelism such as data and tensor parallelism and memory efficient blockwise attention and feedforward (Rabe and Staats, 2021; Dao et al., 2022; Liu and Abbeel, 2023b), which we will show in experiment Section 5.
@@ 5.1 Evaluating Max Context Size
We evaluate maximum supported context length using fully sharded tensor parallelsim (FSDP) (Facebook, 2023) which is widely used in prior end-to-end training (Touvron et al., 2023; Geng and Liu, 2023). We note that no tensor parallelism is considered in our evaluations since our approach is independent of tensor parallelism. Practitioners can combine our method with tensor parallelism, which we will show in Section 5.2. Using FSDP allows us to set the same batch size in tokens for baselines and our approach, ensuring a fair comparison. Concretely, on $n$ devices, FSDP is used to shard the model for baselines, which gives a sequence length of $l$. The total batch size in tokens is $nl$. We utilize FSDP along with Ring Attention to extend the sequence length to $\frac{nl}{m}$ and $m$ sequences. This means that the total batch size in tokens remains the same, but Ring Attention enables a significantly larger context size. Table 3 summarizes the results of our experiments.
Our Ring Attention model consistently surpasses baselines, delivering superior scalability across diverse hardware setups. For example, with 32 A100 GPUs, we achieve over 1 million tokens in context size for 7B model, a 32 times improvement over previous best. Furthermore, when utilizing larger accelerators like TPUv4-512, Ring Attention enables a 256 times increase in context size, allows training sequences of over 30 million tokens. Furthermore, our Ring Attention model scales linearly with the number of devices, as demonstrated by the 8x improvement over previous best on 8 A100 and the 256x improvement on TPUv3-512. If a model can be trained with context size $s$ on $n$ GPUs using the blockwise attention and feedforward, with our Ring Attention approach, it becomes possible to train a model with a context size of $ns$.
Table 3: The maximum context length supported in end-to-end training using fully sharded data parallelism and various transformers architectures. We show different model sizes and accelerators. Baselines are vanilla transformer (Vaswani et al., 2017), transformer with memory efficient attention (Rabe and Staats, 2021), and transformer with memory efficient attention and feedforward (Liu and Abbeel, 2023b). The context size is reported in tokens (1e3). Our Ring Attention substantially outperforms baselines and enables training sequences that are up to device count times longer than prior state-of-the-arts.
@@
7B 4 64 128 4096 32x
13B 4 32 64 2048 32x
TPUv3-512
7B 1 4 8 2048 256x
13B 1 2 8 1024 128x
TPUv4-1024
3B 8 16 32 16384 512x
7B 4 8 16 8192 512x
13B 4 8 16 4096 256x
30B 2 4 8 2048 256x
@@
3B 4 8 32 4096 128x
7B 2 8 16 2048 128x
Table 4: Model flops utilization (MFU) with different training configurations: model sizes, compute, and context lengths. Ring Attention enables training large models (7B-65B) on large input context sizes (over 4M) with negligible overheads.
Model size 7B 13B 13B 30B 65B
Compute 8x A100 8x A100 32x A100 TPUv4-1024 TPUv4-1024
Memory efficient attention & FFN Context size (×$\times 1\mathrm{e}3$) 32 16 64 16 8
Ring Attention Context size (×$\times 1\mathrm{e}3$) 256 128 2048 2048 1024
@@ 5.2 Evaluating Model Flops Utilization
We evaluate the model flops utilization (MFU) of Ring Attention in standard training settings using fully sharded data parallelism(FSDP) (Facebook, 2023) and tensor parallelism following LLaMA and OpenLLaMA (Touvron et al., 2023; Geng and Liu, 2023) with Jax SPMD. The batch size in tokens are 2M on 8/32x A100 and 4M on TPUv4-256. Our goal is investigating the impact of model size and context length on MFU, a critical performance metrics while highlighting the benefits of our approach. Table 4 presents the results of our experiments on MFU for different model sizes and context lengths. We present the achieved MFU using state-of-the-art memory efficient transformers BPT (Liu and Abbeel, 2023b), compare it to our anticipated MFU based on these results, and demonstrate the actual MFU obtained with our approach (Ring Attention). For fair comparison, both BPT and our approach are based on the same BPT implementation on both GPUs and TPUs.